# Accurate BGV Parameters Selection: Accounting for Secret and Public Key Dependencies in Average-Case Analysis

### v2

@@ 4.3 Homomorphic multiplications
Theorem 4.1 provides an upper bound on the variance of the product $a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}}$. However, it is significant to point out that this estimate can be regarded as a faithful approximation, i.e.,
$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})\approx n\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}})|_{i})\mathsf{Var}((a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}).$
The reason behind this statement can be immediately found in the distributions of the $b_{\mu}$-terms in the noise of a fresh ciphertext. In fact, for a fresh ciphertext, the critical quantity can be expressed as
$\nu=a_{0}+a_{1}s=(b_{0}(0)+b_{1}(0)\cdot e)+b_{0}(1)s,$
where
Then, it can be observed that the variance of the $b_{\mu}$-values is given by
$\begin{cases}\mathsf{Var}(b_{0}(0)|_{i})=V_{m}+t^{2}V_{e}\\ \mathsf{Var}(b_{1}(0)|_{i})=nt^{2}V_{e}V_{u}\\ \mathsf{Var}(b_{0}(1)|_{i})=nt^{2}V_{e}V_{s}\end{cases}$
where $n$ typically assumes values in $\{2^{13},2^{14},2^{15},2^{16}\}$. Consequently,
$\mathsf{Var}(b_{0}(0)|_{i})\ll\mathsf{Var}(b_{1}(0)|_{i}),\mathsf{Var}(b_{0}(1)|_{i}).$
Therefore, $b_{0}(0)$ becomes negligible in our analysis, and we can assume $a_{0}\approx b_{1}(0)\cdot e$, obtaining the approximated version of the theorem.
However, in order to maintain our results as general as possible, our next analysis will refer to the upper bound version of the theorem, still knowing that it corresponds to a reliable approximation.
5 Application of Our Variance Estimation Method to Fixed Circuits
This section employs the results of Section 4 to propose an effective method for tracking the error growth in circuits with a fixed number of operation. As previously highlighted, obtaining bounds sufficiently tight with respect to the experimental values of the error coefficients’ variance is crucial for development and real use of homomorphic encryption.
These bounds, in fact, allow one to construct new techniques for efficiently finding initial scheme parameters that improve performance and security simultaneously. Naturally, the behavior of the error is very dependent on the specific circuit under consideration, that is, the sequence of operations that the scheme must accommodate.
We analyze a circuit in which pairs of ciphertexts are progressively multiplied. We focus on circuits that primarily involves homomorphic multiplication, since, as previously mentioned, it is the most complex and therefore the most significant operation to study.
For clarity, we divide the analysis into two distinct scenarios. In the first scenario, only theoretical, the modulus switch technique is not applied, resulting in a very rapid increase of the error. Instead, in the second case, the modulus switch is performed after each multiplication.
Typically, the first scenario is not taken into account, as the error magnitude, without modulus switching, quickly becomes unmanageable, making the scheme impractical for real-world applications. In contrast, with modulus switching, error growth is significantly diminished, and the noise remains within a manageable range. This is, in fact, one of the primary goals pursued in designing practical FHE schemes.
However, the application of modulus switch drastically reduces the differences in estimates derived from even significantly distant approaches. Instead, by studying circuits without the modulus switch, we believe that the necessity of considering the dependencies in the coefficients of the error polynomial becomes unequivocal.
For this reason, although the absence of modulus switching does not reflect a realistic scenario, we decided to start our analysis in this simplified context. This allows us to highlight the key structural aspects of error growth, which are then carried over and extended to the more realistic case involving modulus switching.
Again, we want to stress that circuits without modulo switch are not practicable and do not have real world applications. The choice to describe this scenario is purely to complete the explanation of our method, further emphasizing the differences between our approach and others considered.
The remainder of this section is structured as follows. In the first part, we present circuits without modulus switching. Subsequently, we focus on practical circuits that employ modulus switching, showing how the results from Section 4 can be applied in this setting.
Based on the results, we are confident that our estimates represent a tangible improvement over the current state of the art. Our bounds closely match the experimental results, while consistently avoiding underestimation, which affects instead the existing approaches and must be avoided for the proper functioning of the scheme together with its security.
5.1 Circuits without Modulo Switch
Figure 1: Reference circuit
Figure 1 provides a schematic representation of the circuit used as a reference for our analysis. Let $M$ denote the multiplicative depth of the circuit, that is, the number of multiplications to be supported. To facilitate the analysis, we divide the circuit into $L=M+1$ levels.
For the considered circuit, we assume the independent and random generation of $2^{L-1}$ plaintexts, which are then encrypted using the same public key, resulting in $2^{L-1}$ fresh ciphertexts, denoted by $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$.
At each level $l\in[0,\dots,L-1]$, pairs of ciphertexts are multiplied, yielding the corresponding ciphertexts of the $(l+1)$-th level, and then the process is repeated until the final ciphertext is computed.
Our study aims to estimate the variance of the coefficients of the critical quantity $\nu^{(l)}|_{i}$, denoted as $V_{l}$, which corresponds to the ciphertexts of the circuit at level $l$.
According to the current average-case approach, the variance at level $l$ can be approximated as
$V_{l}\leq nV_{l-1}^{2}.$
However, as pointed out by the authors of [25], this estimation tends to underestimate the variance observed in experimental evaluations. Such underestimation may lead to security vulnerabilities, as previously discussed, and, therefore, must be avoided.
In contrast, in our investigation, we proved that
$V_{l}\leq nh(l)V_{l-1}^{2},$ (5)
where $h$ is a function specifically designed to account for the dependencies among the coefficients of the error polynomial. This construction is based on the correction function $F$ introduced in [2] and represents the crucial difference compared to the current approach [25].
Section 4 has laid the groundwork for deriving this bound. Thus, the construction of our bound can now be presented, beginning with the initial levels and subsequently extending the reasoning to the general case for a level $l$.
First Multiplication
Let $\bm{c},\bm{c^{\prime}}$ be two fresh ciphertexts with critical quantities $\nu=a_{0}+a_{1}s$, $\nu^{\prime}=a^{\prime}_{0}+a^{\prime}_{1}s$, respectively, where
$\begin{cases}a^{\prime}_{0}=b^{\prime}_{0}(0)+b^{\prime}_{1}(0)\cdot e=(m^{\prime}+te^{\prime}_{0})+tu^{\prime}\cdot e\\ a^{\prime}_{1}=b^{\prime}_{0}(1)=te^{\prime}_{1}\end{cases}$
Let $V_{0}$ be the variance of the critical quantity associated with the fresh ciphertexts. Then, it can be observed that
$\begin{cases}V(a_{0}|_{i})=V(a^{\prime}_{0}|_{i})=t^{2}(\frac{1}{12}+V_{e}+nV_{e}V_{u})\\ V(a_{1}s|_{i})=V(a^{\prime}_{1}s|_{i})=t^{2}nV_{e}V_{s}\end{cases}$
Therefore, it is possible to assume that
$V(a_{0}|_{i})\approx V(a_{1}s|_{i})\approx\frac{V_{0}}{2},$ (6)
since $V_{u}=V_{s}$, due to the choice of the distributions of $s$ and $u$, and
$\frac{1}{12}+V_{e}\ll nV_{e}V_{u},$
which is then negligible, as highlighted in 1.
Clearly, the same reasoning can be applied to $V(a^{\prime}_{0}|_{i}),V(a^{\prime}_{1}s|_{i})$, obtaining
$V(a^{\prime}_{0}|_{i})\approx V(a^{\prime}_{1}s|_{i})\approx\frac{V_{0}}{2}.$ (7)
Our goal is to estimate the coefficients variance of $\nu_{\mathsf{mul}}=\nu\cdot\nu^{\prime}$, which will be referred to as $V_{1}$, following the level notation introduced at the beginning. Moreover, according to the previously introduced notation, it is possible to write $\nu_{\mathsf{mul}}$ as
$\nu_{\mathsf{mul}}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a_{1}a^{\prime}_{0})\cdot s+a_{1}a^{\prime}_{1}\cdot s^{2}.$
Consequently, when the variance is considered, applying Theorem 4.1 and using that the covariance of this addends is zero, it follows that
$V_{1}:=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$ $\leq n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(0,0)}{F_{e}(1,1)}$
$+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(0,1)}{F_{e}(1,0)}$
$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(1,0)}{F_{e}(0,1)}$
$+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}{F_{e}(0,0)}.$
Moreover, this bound can be simplified, observing that $F_{s}(0,0)=F_{s}(1,0)=F_{s}(0,1)=1$ and $F_{e}(0,0)=F_{e}(1,0)=F_{e}(0,1)=1$. Therefore, it follows that
$V_{1}\leq$ $\,n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{e}(1,1)}+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i})$
$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i})+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}.$
Again, this can be reformulated using (6) and (7) as
$V_{1}$ $\leq\frac{nV_{0}^{2}}{4}\left(F_{e}(1,1)+{F_{s}(1,1)}+2\right).$
Hence, the first value of the function $h$ we aim to construct, can be fixed as
$h(1)=\frac{F_{s}(1,1)+{F_{e}(1,1)}+2}{4}.$
Thus, for the first level, the following inequality has been proven
$V_{1}\leq nh(1)V_{0}^{2}.$
Second Multiplication
In the second multiplication, the ciphertexts involved will have an associated critical quantity of the form
$\nu=a_{0}+a_{1}\cdot s+a_{2}\cdot s^{2},$
where, according to our study, each coefficient has variance
$\begin{cases}\mathsf{Var}(a_{0}|_{i})\approx\frac{F_{e}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{1}s|_{i})\approx\frac{2V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{2}s^{2}|_{i})\approx\frac{F_{s}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\end{cases}$
Therefore, by repeating the same procedure depicted above, the critical quantity after the second multiplication will be of the form
$\nu\cdot\nu^{\prime}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a^{\prime}_{0}a_{1})s+(a_{1}a^{\prime}_{1}+a_{0}a^{\prime}_{2}+a^{\prime}_{0}a_{2})s^{2}+(a_{1}a^{\prime}_{2}+a^{\prime}_{1}a_{2})s^{3}+a_{2}a^{\prime}_{2}s^{4}$
Thus, its coefficients’ variance can be estimated as
$V_{2}$ $\leq\frac{nV_{1}^{2}}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}$
$\quad+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]},$
where this inequality follows from the fact that $a_{0}$ contains a term multiplied by $e^{2}$, $a_{1}$ contains a term multiplied by $e$ and $a_{2}$ does not contain any powers of $e$. As a result, $h(2)$ can be chosen as
$h(2)$ $=\frac{1}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}$
$\quad+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]}.$
Consequently, $V_{2}\leq nh(2)V_{1}^{2}$.
Generalization
Finally, this procedure can be easily generalized in order to compute all the necessary values for $h$, therefore proving our bound in (5), i.e.,
$V_{l}\leq nh(l)V_{l-1}^{2},\hskip 28.45274ptl=1,\dots,L-1.$
5.2 Circuits with Modulo Switch
Figure 2: Reference circuit
In the present section, the reasoning previously introduced is adapted to circuits where the modulus switching technique is employed, in order to exploit the advantages of the method in practical scenarios.
Let $L$ denote the number of levels in the circuit, and $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$ the initial fresh ciphertexts, generated by encrypting $2^{L-1}$ independent and randomly generated messages, using the same key. The reference circuit for our analysis is depicted in Fig. 2.
We consider a model in which, at each level, homomorphic multiplication of pairs of ciphertexts is carried out. The key difference is that at the beginning of each level $l\in\{1,\dots,L-1\}$ the input ciphertexts are switched to a smaller modulus, to maintain the error almost constant throughout the process. In particular, following the previously defined notion, let $q_{l}$ be the ciphertexts moduli of each level, defined as
$q_{l}=\prod_{j=0}^{l}p_{j},$
where $p_{j}$ are primes such that $\gcd(p_{i},p_{j})=1$ for $i\neq j$.
The process represented by the circuit can be resumed as follows:
  • At level 0, all the fresh ciphertexts are defined in $\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$;
  • At level 1, modulo switch to $q_{L-2}$ is applied to each of the initial fresh ciphertexts. After that, the first homomorphic multiplication is performed in $\mathcal{R}_{q_{L-2}}$, yielding $2^{L-2}$ resulting ciphertexts in $\mathcal{R}_{q_{L-2}}^{3}$. Finally, these ciphertexts are relinearized to obtain equivalent ones in $\mathcal{R}_{q_{L-2}}^{2}$, which constitutes the input of the next level;
  • At level $2$, each input ciphertext undergoes a modulo switch from $q_{L-2}$ to $q_{L-3}$, which is again followed by the homomorphic multiplication of pairs of them. To conclude, a relinearization step is carried out;
  • This process is repeated until level $L-1$, where a final ciphertext, which constitutes the output of the circuit, is returned;
It should be noted that, in the circuit represented in Fig. 2, the relinearization step is not explicated. In fact, as for the case of the circuit without modulus switching, we decided to omit its contribution from our error analysis, since it is negligible compared to the impact of modulus switch and multiplication.
Before discussing the results, it should be noted that, in contrast to the case without the modulus switching, in this scenario, $L$ ciphertext moduli are required for the operation of the circuit. Therefore, the selection of parameters will consist in choosing all the $q_{l}$ values, i.e., all the primes $p_{j}$, such that the correct functioning of the scheme is ensured while still choosing sufficiently small moduli in order to improve efficiency, which is crucial for practical usage of BGV.
For better understanding, we will divide our analysis according to the level $l$ considered. Specifically, an in-depth explanation of the noise estimates for the first three levels will be provided. Then, these results will be generalized for a fixed level $l$, thus delineating our bound.
Level 0
Let $\bm{c}\in\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$ be a fresh ciphertext, obtained encrypting a random message $m\in\mathcal{R}_{t}$.
Let $\nu_{\mathsf{clean}}=a_{0}+a_{1}s$ be the critical quantity associated with $\bm{c}$.
As for the circuits without modulo switching, it is possible to assume that
$\mathsf{Var}(a_{0}|_{i})\approx\mathsf{Var}((a_{1}s)|_{i})\approx\frac{V_{0}}{2},$ (8)
where the notation $V_{0}$ is used to denote the variance of the coefficients of the critical quantity $\nu_{\mathsf{clean}}$. In general, the notation $V_{l}$ will be used to represent the variance of the coefficients of the critical quantity at the end of level $l$. It is significant to observe that the assumption in (8) applies to all initial $2^{L-1}$ ciphertexts. In fact, the estimate of $V_{l}$ is the same for all ciphertexts belonging to the same level.
Level 1
As mentioned above, at the beginning of level $1$, all ciphertexts are subject to a modulus switch from $q_{L-1}$ to $q_{L-2}$.
Therefore, the error and the modulus of $\bm{c}$ are rescaled by a factor of ${q_{L-2}}/{q_{L-1}}={1}/{p_{L-1}}$. Thus, the critical quantity associated to the resulting ciphertext is given by
$\nu_{\mathsf{modswitch}}=\frac{\nu_{\mathsf{clean}}+\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}},$ (9)
where $\nu_{\mathsf{ms}}(p_{L-1})$ is defined as
$\nu_{\mathsf{ms}}(p_{L-1})=\delta_{0}+\delta_{1}s\hskip 28.45274pt\delta_{i}=t[-c_{i}t^{-1}]_{p_{L-1}}$
In particular, for the following analysis we will use the notation
$\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}=a_{0}^{\mathsf{ms}}+a_{1}^{\mathsf{ms}}s,$
with
$\begin{cases}a_{0}^{\mathsf{ms}}=\frac{\delta_{0}}{p_{L-1}}\\ a_{1}^{\mathsf{ms}}=\frac{\delta_{1}}{p_{L-1}}\end{cases}$
Moreover, we will denote with $V_{\mathsf{ms}}$ the following variance
$V_{\mathsf{ms}}=\mathsf{Var}\left(\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}\middle|_{i}\right).$
For our analysis, it is possible to assume that $V_{\mathsf{ms}}\approx\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$. This is a consequence of the fact that, observing the distributions of the coefficients in $a_{0}^{\mathsf{ms}}$ and $a_{1}^{\mathsf{ms}}$, it is evident that $\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})\ll\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$. In fact, noting that $c_{i}\leftarrow\mathcal{U}_{q_{L-1}}$, it is possible to estimate the variance of $a_{0}^{\mathsf{ms}}|_{i}$ and $a_{1}^{\mathsf{ms}}|_{i}$ as
$\begin{cases}\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})=\mathsf{Var}(a_{1}^{\mathsf{ms}}|_{i})=t^{2}/12\\ \mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})={t^{2}nV_{s}}/{12},\end{cases}$
where $n$ is typically in $\small\{2^{12},2^{13},2^{14},2^{15}\}$. With this in mind, we can proceed with the analysis of the noise growth related to the multiplication of two ciphertexts, whose associated critical quantity is as in (9).
We start by observing that the critical quantity in (9) can be written as
$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+a_{0}^{\mathsf{ms}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
Moreover, it is possible to further simplify the notation by using the fact that $\mathbb{E}[a_{0}^{\mathsf{ms}}]=0$ and its variance is negligible. Therefore, we rewrite this value as
$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
Then, it can be derived that
$\begin{cases}\mathsf{Var}(\hat{a_{0}}|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}\\ \mathsf{Var}(\hat{a_{1}}s|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}+V_{\mathsf{ms}}\end{cases}$
since
In order to estimate the variance of the critical quantity resulting from multiplication, we will rely on Theorem 4.1. In particular, in this analysis, we will try to avoid explicitly writing the terms $b_{\mu}(\iota)$ involved in the expression of a generic $a_{\iota}=\sum_{\mu=0}^{K}b_{\mu}(\iota)e^{\mu}$, to avoid the notation becoming cumbersome. However, it is crucial to keep in mind the highest power of $e$, namely $K$, present in the terms $a_{\iota}$ involved in the multiplications, as the correction of the function $F$ depends on this value. Precisely, for fresh ciphertexts at level 00, it can be seen that for $a_{0}$ the associated $K$ is $1$, while for $a_{1}$ we have $K=0$. Instead, for the terms $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$ the case is quite different. In fact, these values can be assumed to be randomly distributed over $\mathcal{R}_{t}$, i.e. $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}\leftarrow\mathcal{U}_{t}$ for every level of the circuit. This makes the analysis slightly more complicated, since, in order to derive tight estimates, it will be necessary to distinguish the contribution of ${a_{0},a_{1}}$ from the one of $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$.
Now, assume that $\bm{c},\bm{c}^{\prime}\in\mathcal{R}_{q_{L-2}}\times\mathcal{R}_{q_{L-2}}$ are two ciphertexts at level 1, after modulus switching has been carried out. Their associated critical quantity is given, respectively, by
$\begin{cases}\nu_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a_{0}+\left(\frac{1}{p_{L-1}}a_{1}+a_{1}^{\mathsf{ms}}\right)s\\ \nu^{\prime}_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a^{\prime}_{0}+\left(\frac{1}{p_{L-1}}a^{\prime}_{1}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s\\ \end{cases}$
$\nu_{\mathsf{mul}}^{(1)}=a_{0}^{\mathsf{mul}}+a_{1}^{\mathsf{mul}}s+a_{2}^{\mathsf{mul}}s^{2},$
namely,
$\nu_{\mathsf{mul}}^{(1)}$ $=\frac{1}{p_{L-1}^{2}}a_{0}a^{\prime}_{0}+\frac{1}{p^{2}_{L-1}}\left(a^{\prime}_{0}{a_{1}}+a_{0}{a^{\prime}_{1}}\right)s+\frac{1}{p_{L-1}}(a^{\prime}_{0}a_{1}^{\mathsf{ms}}+a_{0}{a^{\prime}_{1}}^{\mathsf{ms}})s\hskip 4.26773pt$
$\hskip 11.38092pt+\frac{1}{p^{2}_{L-1}}a_{1}a^{\prime}_{1}s^{2}+\frac{1}{p_{L-1}}(a_{1}{a^{\prime}_{1}}^{\mathsf{ms}}+a^{\prime}_{1}{a_{1}}^{\mathsf{ms}})s^{2}+{a^{\prime}_{1}}^{\mathsf{ms}}{a_{1}}^{\mathsf{ms}}s^{2}.$
So, we are now able to provide an estimate of the variance $V_{1}=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$, applying Theorem 4.1, as follows
$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{0}|_{i})F_{e}(1,1)$
$+\frac{n}{p_{L-1}^{4}}V(a^{\prime}_{0}|_{i})V(a_{1}s|_{i})+\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{1}s|_{i})$
$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{0}|_{i})V(a_{1}^{\mathsf{ms}}s|_{i})+\frac{n}{p_{L-1}^{2}}V(a_{0}|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})$
$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})V(a^{\prime}_{1}s|_{i})F_{s}(1,1)+\frac{n}{p_{L-1}^{2}}V(a_{1}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)$
$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{1}s|_{i})V({a_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)+{n}V({a_{1}}^{\mathsf{ms}}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1).$
which, recalling that $\operatorname{Var}(a_{\iota}|_{i})=\operatorname{Var}(a^{\prime}_{\iota}|_{i})$ and that $V({a_{1}}^{\mathsf{ms}}s|_{i})=V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})=V_{\mathsf{ms}}$, can be rewritten as
$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})^{2}F_{e}(1,1)+$
$+\frac{2n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a_{1}s|_{i})+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{0}|_{i})$
$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})^{2}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{1}s|_{i})+nF_{s}(1,1)V_{\mathsf{ms}}^{2}.$
By substituting to $V(a_{0}|_{i})$,$V(a_{1}s|_{i})$ the value $\frac{V_{0}}{2}$, our bound on $V_{1}$ is derived. To conclude, it is possible to observe that the following estimate holds
[…]

### v3

@@ 6 Variance Estimation in Circuits with Fixed Depth
At first glance, it might seem that, in order to correctly estimate the variance of the multiplication, it is sufficient to apply a factor of 2, arising from the correction of the product between two $s$-terms, thus making it unclear why the dependence on the public key should be considered. It is crucial to emphasize that these estimates are strictly related to a circuit model that we are specifically constructing to guarantee the Gaussianity of the error coefficients according to Equation 12. However, if one wanted to accurately study the variance in different circuits, where the term $V_{\mathsf{ms}}$ is not the only significant contribution, the need to account for these dependencies would become even more evident. This necessity reaches its maximum importance when considering circuits in which modulus switching is not applied. Since the use of modulus switching is mandatory for the practical implementation of BGV, in this work we omitted the results obtained from the analysis of such circuits. However, whether modulus switching is not used, or it is applied with primes that do not satisfy Equation 12, it remains crucial to take into account all dependencies arising from both $e$ and $s$ as well as from their powers. As an example of the applicability of our approach, if error coefficient distributions that are non-Gaussian, but still suitable for use, were considered, the variance analysis we propose would remain valid and could still provide accurate estimates that never underestimate the experimental values (see Theorem 4.2).
7 Comparison with Previous Works
When estimating the error in schemes such as BGV, existing approaches can broadly be divided into two main categories: worst-case and average-case. The former includes analyses based on the Euclidean norm, as in [6], the infinity norm, as in [22, 30], and the canonical norm, as in [15, 17, 25, 29, 34], with the latter providing the tightest bounds among the worst-case analyses. In contrast, average-case approaches model the noise coefficients as random variables and focus on their mean and variance to derive probabilistic bounds on the minimum ciphertext modulus required to guarantee correctness.
Although the latter approaches appear more promising in reducing the ciphertext modulus size, research for a general method to estimate the variance remains incomplete. Most works, such as [35], treat the error coefficients as independent, which, as pointed out in [4, 16, 35], leads to imprecise bounds and often underestimates the modulus size required to prevent decryption failures. Consequently, the scheme may be exposed to potential vulnerabilities, including key-recovery attacks [10]. Other types of methods, by contrast, avoid such underestimations or failures but are limited to specific settings, as in [18], where the authors analyze the error for HElib [28].
The objective of our work is to develop an average case approach that takes into account the intrinsic dependencies introduced by the secret and public key, allowing to obtain variance estimates very close to experimental values, without ever underestimating them. At the same time, our aim is to propose a method independent of the specific library or circuit considered, providing the theoretical basis necessary to build ad hoc estimates depending on the circuit and library of interest. Although, to highlight the effectiveness of our approach, we have focused on circuits that exclusively include multiplications preceded by the modulus switch, our analyses and estimates remain valid for any type of homomorphic circuit. Our method is in fact intrinsically modular: each homomorphic operation is analyzed independently, producing variance bounds that can be composed to precisely estimate the overall noise growth of a given circuit, assuming only that the multiplications are preceded by a modulus switch – a completely reasonable hypothesis in the context of BGV typical usage.
In this section, we demonstrate the efficacy of our average-case approach by comparing it to the state-of-the-art works. Specifically, we illustrate how our average-case analysis provides tighter and more practical bounds than traditional worst-case methods [34], and we further compare it with other average-case approaches [35, 18]. We then show how our choice of ciphertext moduli achieves a substantial size reduction compared to those adopted in widely used libraries such as OpenFHE [1] and HElib [28].
Canonical norm.
For the comparison with the worst-case approach, we specifically refer to the estimates proposed in [34].
   Homomorphic operation    Error bounds with canonical norm
   Enc    $\nu_{\mathsf{clean}}||^{\text{can}}\leq Dt\sqrt{n\left(1/12+2nV_{e}V_{s}+V_{e}\right)}$
[…]
Table 2: Canonical norm depending on the homomorphic operations.
[…]
In contrast, average-case approaches, such as the one proposed in this work, allow significantly tighter bounds by imposing the condition $\|\nu\|_{\infty}\leq D\sqrt{2V}$, where $V$ denotes the variance of each coefficient of $\nu$. According to Equation 2, this bound holds with probability at least $1-n(1-\operatorname{erf}(D))$, which, for $D=6$, exceeds $1-2^{-40}$. However, in practical scenarios, choosing $D=8$ is preferable, as it limits the failure probability to $2^{-77}$, when $n\leq 2^{15}$. Note that this bound better captures the actual distribution of the error and results in much tighter predictions after multiplications. In fact, by Section 6, after a multiplication that follows modulo switch, we have $\nu_{\text{mul}}||_{\infty}\leq 2D\sqrt{n}V_{\mathsf{ms}}$.
Current Average-Case Approaches.
Within average-case analysis frameworks, the main novelty of our approach lies in tracking the dependencies between the error coefficients, particularly those arising from the multiplication of two ciphertexts. Specifically, given the critical quantities $\nu$ and $\nu^{\prime}$ of two ciphertexts with coefficient variances $V$ and $V^{\prime}$, respectively, the current average-case methodology [35] yields the expression $\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx nVV^{\prime}$. Therefore, assuming that two ciphertexts have just undergone modulus switching, the variance resulting from their product is, in accordance with the previously established considerations
$\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx nV_{\mathsf{ms}}^{2}\,.$
In contrast, our method estimates this value according to Section 6 , resulting in
$\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx 2nV_{\mathsf{ms}}^{2}\,.$
The factor 2, which arises from accounting for the dependencies among the critical quantities, allows for extremely accurate values, as shown in Table 1.
From this comparison, it should be evident that accounting for the dependencies in the coefficients of the error polynomial is fundamental for obtaining accurate and correct estimates of the experimental variance.
In particular, we believe that this consideration is precisely what overcomes the underestimation inherent in the approach presented in [35].
It is crucial to point out that the factor of 2 is specific to the type of circuit we constructed, in which the variance is independent of the circuit level. However, when analyzing errors in circuits where this condition is not met, it is still possible to derive general upper bounds using Theorem 4.2, once again obtaining estimates that never underestimate the error. Nevertheless, for the reasons outlined above, including guaranteeing the Gaussianity of the error, we recommend choosing primes that allow the variance of the error after the modulus switch to be well approximated by $V_{\mathsf{ms}}$, as explained previously.
Our results appear to be very close to those reported in [18]. In particular, it can be observed that a factor of 2 also appears in [18, Lemma 9].
Current Libraries.
Finally we compare the ciphertext modulus $q$ as estimated by our method against those adopted by two of the most widely used libraries: OpenFHE [1] and HElib [28]. We recall that the ciphertext modulus must be selected to ensure that $\|\nu\|<\frac{q}{2}$. In our selection, we additionally require (12) to guarantee the Gaussian behavior of the error.
OpenFHE.
In OpenFHE [1], the generation of the moduli depends on the multiplicative depth selected as input, in a manner similar to the choice of moduli described in this paper. Moreover, the way multiplications are executed in a fixed-depth circuit of the library is consistent with the circuits analyzed in our work, where each multiplication between two ciphertexts is preceded by a modulus switching operation. For a circuit of multiplicative depth $M$, $M+2$ moduli are generated, to allow for an additional modulus switching after the final multiplication. Therefore, we generate the moduli in a way consistent with the library’s design, to allow a fair comparison between the modulus sizes used in OpenFHE and those proposed in this paper.
In Figure 5, we compare the ciphertext modulus sizes $q$ for circuits with multiplicative depths 3 (Figure 5(a)) and 6 (Figure 5(b)). The values are obtained using our method (denoted as our) and those generated by OpenFHE (denoted as OpenFHE). Note that, for the generation of our parameters, we fixed $D=8$.
[…]